OpenStudy (anonymous):

Assuming biological substances are 90 percent water and the density of water is 1.0 x 10^3 kg/m^3, estimate the masses (density multiplied by volume) of the following. a. a spherical cell with a diameter of 1.0 um (volume = 4/3 pi r^3 b. a fly, which can be approximated by a cylinder 4.0 mm long and 2.0 mm in diameter (volume = lpir^2

OpenStudy (anonymous):

what do u mean by 1.0 um...

OpenStudy (anonymous):

It means 10^-6 m like how cm is 10^-2 m

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

mass of water in (a) mass of water = 1000kg/m^3 radius of sphere = 0.5 x 10^-6 m \[V = \frac{ 4 }{ 3 } \pi r^3 \] substituting, values of r, then you'll get 5.236 x 10^-19 m^3 Hence weight of water in that sphere is 5.236 x 10^-19 x 1000 = 5.236 x 10^-16 kg

OpenStudy (anonymous):

hold on.. solving b

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

same goes.. radius is 10^-3 m \[V= \pi r^2 l \] then substituting values of r and l then you'll get 1.2566 x 10^-8 m^3 multiplying with 1000 then you get 1.2566 x 10^-5 kg. or can be simplified to 12.57 microgram (correct to 4 s.f.)

OpenStudy (anonymous):

solved

OpenStudy (anonymous):

ok let me check it really fast. Thanks for the help

OpenStudy (anonymous):

I got 12566 for some reason

OpenStudy (anonymous):

thats 5 signifiant figures. mine i rounded it to 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

also do you mind helping me out with one more problem hahaha :)

OpenStudy (anonymous):

ok o.o btw vote me best response xD

OpenStudy (anonymous):

I WILL BE HAPPY TO DO IT AGAIN :)

OpenStudy (anonymous):

haha posted your qst?

OpenStudy (anonymous):

So what is the units for both of the answers. Kg or m^3 also I got 1.25 x 10^-2 kg

OpenStudy (anonymous):

I got 1.25 times 10^-2 mkg

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hmm my aswer here i recalculated still its 1.2566 x 10^-5 kg

OpenStudy (anonymous):

in standard form kg

OpenStudy (anonymous):

thats weird

OpenStudy (anonymous):

I did pi (0.002/2)^2 (4)

OpenStudy (anonymous):

i just did diameter over 2 instead of using radius

OpenStudy (anonymous):

I do get 1.257 x 10^-5 for m^2 for volume?

OpenStudy (anonymous):

u cant use diameter, you need to use radius...

OpenStudy (anonymous):

ok just solve it step by step and show me what you did

OpenStudy (anonymous):

then I sware to let you go :)

OpenStudy (anonymous):

radius is 10^-3 m length 4.0 x10^-3 m V = pi (10^-3)^2 (4.0 x10^-3) = 1.2566 x 10^-8 multiply 1000 = 1.2566 x 10^-5 kg

OpenStudy (anonymous):

i forgot the length. Now i will try it out lol

OpenStudy (anonymous):

lol......

OpenStudy (anonymous):

I get the same answer but isn't it meters(kg) in units

OpenStudy (anonymous):

?

OpenStudy (anonymous):

lol i gtg d sry ><

OpenStudy (anonymous):

can you look into it please?